OpenStudy (anonymous):

*WILL GIVE MEDAL* Can someone help me find the parabola of f(x)=2(x-6)^2+4

OpenStudy (anonymous):

I already have it graphed, I think I need the points.

OpenStudy (anonymous):

f(x)=2(x-6)^2+4 mix at f'(x) = 0 f'(x) =4(x-6) = 0, so x = 6 and y = f(6) = 4 thus the minimum is (6,4)

OpenStudy (anonymous):

what other points do you need?

OpenStudy (anonymous):

oh, and mix is a typo, I meant min

OpenStudy (anonymous):

@Sherabartlett is it just the minimum? Or do you need other points as well?